OpenStudy (anonymous):

Solve: c^2-4c=30+3c With answer please, checking my work.

OpenStudy (accessdenied):

You would move all of the terms to one side leaving the other side " = 0". Then, you could check if it factors or otherwise use quadratic equation to solve... What were the answers that you got?

OpenStudy (anonymous):

Not the right answers loool, totally forgot to move the terms. Thanks.

OpenStudy (accessdenied):

No problem. :) c^2 - 7c - 30 = 0 (c - 10)(c + 3) = 0 c - 10 = 0 c = 10 c + 3 = 0 c = -3